OpenStudy (anonymous):

There are n teams in a league and the number of games played is given by the polynomial n^2-n-n. A women's softball league has 13 teams. What is the total number of games to be played?

OpenStudy (amistre64):

is that a typo? -n-n?

OpenStudy (anonymous):

n^2-n=n sry

OpenStudy (amistre64):

nope; still dont make sense

OpenStudy (anonymous):

how about n^2-n=N that is how it is written

OpenStudy (anonymous):

boy, you two are going to pull an all nighter

OpenStudy (amistre64):

ack...nope

OpenStudy (amistre64):

i did 14^2 :)

OpenStudy (amistre64):

169 -13 ---- 156 = N

OpenStudy (anonymous):

so is there a way to write that with varibles

OpenStudy (amistre64):

hmm.... not as written no

OpenStudy (amistre64):

n=13; so 13^2-13 = N if i read it right

OpenStudy (anonymous):

well when I write N=156 it say incorrect

OpenStudy (anonymous):

ok so I will try that

OpenStudy (amistre64):

then theres something your keeping secret from the problem ;)

OpenStudy (anonymous):

the correct answer was 156 by itself w/o the n.....gezzz

OpenStudy (amistre64):

:) picky programs